The ledger of Sarasota Corp. on March 31 of the current year includes the following selected accounts before quarterly adjusting entries have been prepared:

Debit Credit

Prepaid Insurance

$3,900

Supplies

2,550

Equipment

25,900

FV-OCI Investments

177,000

Accumulated Depreciation—Equipment

$8,640

Notes Payable

29,000

Unearned Rent Revenue

8,050

Rent Revenue

58,900

Interest Expense

-0-

Salaries and Wages Expense

13,600


An analysis of the accounts shows the following:

1. The equipment depreciation is $360 per month.
2. One half of the unearned rent was earned during the quarter.
3. Interest of $435 has accrued on the notes payable.
4. Supplies on hand total $920.
5. Insurance expires at the rate of $325 per month.
6.

The FV-OCI Investments were purchased for $177,000 on March 1. No investments were purchased or sold after that date. The fair value on March 31 was $197,000.

If the notes payable have been outstanding since January 1 of the current year, what is the annual interest rate on the notes payable? (Round answer to 2 decimal places, e.g. 15.25%.) what is the annual rate?

Annual Interest Rate
Answer #1

Calculate annual interest rate

Interest expense = Amount*rate*time

435 = 29000*X*3/12

435 = 7250X

X = 435/7250 = 6%